hallm:

2. Write an equation in which the quadratic expression 2x^2 - 2x - 12 equals 0. Show the expression in factored form and explain what your solutions mean for the equation. Show your work.

dude:

Do you know how to factor?

hallm:

no

hallm:

how do you factor?

jhonyy9:

si bc. every terms contain the 2 you can factorizing out and will get 2(x^2 -x -6) = 0 and now for you can factorizing the x^2 -x-6 so you can getting two numbers with sum equal -1 and product equal -6 suppose a and b will be these two numbers so than you can factorizing the x^2 -x -6 = (x-a)(x+b) than a>b bc. a-b need being equal -1 hope helped